OpenStudy (anonymous):

what is the simplified form of (4x^2-20x+25)/(2x-5)

OpenStudy (mertsj):

Factor the numerator and cancel.

jimthompson5910 (jim_thompson5910):

(4x^2-20x+25)/(2x-5) ((2x-5)(2x-5))/(2x-5) ... Note: this is a perfect square 2x-5

OpenStudy (mertsj):

\[\frac{4x ^{2}-20x+25}{2x-5}=\frac{(2x-5)(2x-5)}{2x-5}=2x-5\]

OpenStudy (anonymous):

so if i was given the same choices as last time, its still not equal to 5/2?

jimthompson5910 (jim_thompson5910):

yes to avoid dividing by zero